# Wave-induced drift in third-order deep–water theory

### v1

@@ Abstract
The goal of this work is to investigate particle motions beneath unidirectional, deep-water waves up to the third-order in nonlinearity. A particular focus is on the approximation known as Stokes drift, and how it relates to the particle kinematics as computed directly from the particle trajectory mapping. The reduced Hamiltonian formulation of Zakharov and Krasitskii serves as a convenient tool to separate the effects of weak nonlinearity, in particular the appearance of bound harmonics and the mutual corrections to the wave frequencies. By numerical integration of the particle trajectory mappings we are able to compute motions and resulting drift for sea-states with one, two and several harmonics. Overall, we find the linear theory yields the largest forward drift close to the surface, followed by second and third order. At greater depth this trend reverses, with linear theory yielding the smallest forward drift. Incorporating difference harmonic terms into the classical Stokes drift formulation is found to yield improved agreement with the drift obtained from nonlinear wave theories, particularly at greater depth.
@@ 1 Introduction
Waves on the surface of water are a continuing source of fascination, as well as posing an important scientific and engineering challenge, with impacts on such diverse areas as offshore renewables and marine ecology. While we can observe the myriad forms of waves on the water surface, these shapes are only manifestations of what is going on below the surface: a transport of energy and, to a lesser extent, the movement of water particles themselves. Leonardo da Vinci already observed (Mossa, 2021) that the wave outruns the water, and early mathematical analyses of wave kinematics seemed to confirm this idea. Work in the early 19th century by Green showed that particle paths in linear deep water waves are approximately circular – in seeming agreement with the exact, deep-water solution derived by Gerstner – and the corresponding finite depth particle paths were shown to be approximately elliptical by Airy (Craik, 2004).
@@
Nevertheless, much remains unknown from both an experimental and theoretical perspective. Experiments are notoriously difficult, owing to the challenge of comparing a closed flume with the situation in the (open) ocean. Recent years have seen significant progress, from flume experiments on monochromatic wave trains (Umeyama, 2012) to more recent investigations of particle kinematics for groups and complex flows (Bjørnestad et al., 2021; van den Bremer et al., 2019).
The aim of this work is to explore the theoretical particle trajectories – and drift – associated with nonlinear potential flow theories in infinite water depth. While there are recent mathematical results for the full, nonlinear problem showing that no closed particle paths exist in steady, periodic wave trains (Constantin, 2006; Henry, 2006; Okamoto & Shoji, 2012), these results are not quantitative and are difficult to generalise to nonsteady problems. We will take a ‘weakly nonlinear’ approach, and our starting point will be the Stokes drift formulation which arises in linear (1st order) water wave theory. While a rigorous proof that linear, steady, periodic water waves exhibit a forward drift in the direction of wave propagation (Constantin & Villari, 2008; Constantin et al., 2008) was recently found, the approximate results due to Stokes form the basis of most practical implementations of drift in forecasting codes.
Stokes’ original results for a single wave train were generalised to a spectrum of waves by Kenyon (1969), and the resulting formulation has been extensively explored for different spectral shapes (Webb & Fox-Kemper, 2011), for its sensitivity to the high-frequency spectral tail (Lenain & Pizzo, 2020), and new profiles have been developed for practical implementation (Breivik et al., 2016). Many more references can be found in the recent review by van den Bremer & Breivik (2018).
@@
For a successful comparison we need some way to compute the fluid velocity field , which drives the particle kinematics through the particle trajectory mapping We need a methodology that can deal with several Fourier modes, to simulate irregular seas, and which is accurate to a desired order of nonlinearity. Several options exist, including the use of traditional perturbation theory to obtain a solution to the water wave problem, which has been extended to third-order by Madsen & Fuhrman (2012). An alternative – which we employ – is to use the equivalent Hamiltonian expansion, which is due to Zakharov (1968) and for which the third-order solution has been explored by Zhang & Chen (1999) and Gao et al. (2021).
The remainder of this paper is structured as follows: in Section 2 we introduce the water wave problem and corresponding Hamiltonian. We give more details about the expansion that allows us to extract linear, second, and third-order terms in Section 2.1, discuss the neglect of amplitude evolution for our particle kinematics calculations in Section 2.2, and finally show how the free surface and potential are obtained to each order in Section 2.3. In Section 3 we consider monochromatic waves, which is the natural setting to introduce Stokes drift, and look at the consequences of higher-order theories on particle trajectories. Section 4 is devoted to bichromatic waves. These are the simplest type of non-steady wave-field, but because only two harmonics are involved the linear superposition is always periodic. Finally, Section 5 considers wave groups and random waves involving multiple lowest-order harmonics. Conclusions are presented in Section 6. Appendix A adds some comments on the non-uniqueness of the monochromatic Stokes wave solution at third-order, while Appendix B gives expressions for some coefficients appearing in Section 2.3.
@@ 2 The water wave problem and Hamiltonian formulation
The inviscid, irrotational water wave problem, written in terms of a fluid velocity potential and free surface is
The horizontal () gradient is denoted for clarity. The advantage of this formulation is that all nonlinearities have been moved to the boundary conditions. We also note that, provided the water is deeper than about half a typical wavelength, the still-water depth can be taken to be infinite. This procedure, using
@@
in place of (1d) simplifies calculations considerably. Despite the restriction to irrotational flow and the neglect of compressibility and viscosity, these equations contain a huge range of interesting physics, including many types of wave phenomena.
During the late 1960s, efforts to find a Hamiltonian formulation of the water wave problem (1a)–(1d) came to fruition in the independent work of Zakharov (1968), Miles (1977) and Broer (1974), who found the Hamiltonian
@@ 2.2 Constant magnitude approximation
The procedure outlined above would seem to suggest that we first need to solve the complicated integro-differential Zakharov equation (8). In fact, for certain situations of interest we can circumvent this step. The key insight is that the evolution described by the Zakharov equation occurs on the slow timescale , where is a typical wave steepness and is a typical wave period (e.g. s in the ocean). Therefore, for processes taking place on time-scales faster than 100 wave periods, it may be appropriate to consider the complex amplitudes to be constant.
@@
Incorporating these corrections the complex amplitudes can be written in an approximation up to third-order as
where the small, quadratic terms in the argument reduce to , the well-known Stokes’ correction, when only a single wave-mode is present. This form now contains all the mutual dispersion corrections entailed by the cubically nonlinear water wave theory. It has been demonstrated that taking these corrections into account allows one to predict the short-time propagation of laboratory waves with high accuracy without the need to solve for the evolution of the slow-time variables (Stuhlmeier & Stiassnie, 2021; Galvagno et al., 2021), showing that the constant magnitude approximation is well-justified for short times. Of course, under specific circumstances – for example with a sufficiently narrow spectrum, or close to modulation instability – we may have significant energy transfer (Andrade & Stuhlmeier, 2023). However, given our focus on particle paths (with a characteristic time on the order of one period) the amplitude evolution is negligible (see also the discussion in Stuhlmeier & Stiassnie (2019) and Appendix A therein).
@@ 2.3 Recovery of the third-order solution
The corrections contained in the second-order and third-order free surface elevation essentially change the wave shape. The resulting harmonics of the form behave as waves with wavenumber and frequency but because these sums and differences do not satisfy the dispersion relation they are referred to as “bound waves”, in contrast to the free waves of (11). The well-known impact of these bound waves on sharpening the crests and flattening the troughs is shown in Figure 1 for a bichromatic sea, together with the effect of the third-order frequency correction (12) on the wave phases.
@@ 2.3.1 Recovery of the bulk potential
We relate the bulk potential and the surface trace to one another in terms of an expansion. Taking the problem in infinite depth, we need to solve
@@
3 Particle kinematics for monochromatic waves
We have developed above a general formulation of the third-order water wave problem. However, this formulation in terms of a free surface elevation and a velocity potential needs to be connected to the kinematics of fluid particles. The most transparent starting point is with monochromatic waves – i.e. initial conditions for which the linear theory is described by a single Fourier mode Of course, the lengthy formulation of Section 2.3 is not strictly necessary, since monochromatic solutions to third-order can be readily found in the literature, e.g. Wehausen & Laitone (1960). Consequently, we begin by reviewing some familiar ideas.
3.1 Linear theory
Linear deep-water theory poses the water wave problem on the half-plane below Solving the Laplace equation then yields the well-known solution in deep water
@@ 3.1 Linear theory
It is worth noting that this system has only a tenuous connection to the linear (or weakly nonlinear) theory, but rather represents the link between Lagrangian and Eulerian descriptions of the fluid motion. For example, the idea that the free surface is always composed of the same fluid particles is only nearly satisfied (under the assumption of waves of small amplitude), so that the particle trajectory mapping only nearly describes the dynamics of particles on the surface (or any streamline).
Given that we are dealing with linear theory, it seems simple enough to solve the system (26). It is an unwelcome surprise that inserting the linear potential yields a nonlinear system of differential equations
While this system cannot be solved analytically it can be shown rigorously that its particle trajectories are not closed (Constantin & Villari, 2008).
In lieu of an analytical treatment, two alternatives remain: approximation or numerical solution. The former approach is found in most textbooks on water waves, and begins with a Taylor expansion of the fluid velocity field as
and similarly for Keeping only the lowest order terms (i) in the Taylor expansion yields the explicit system
with which yields the circular particle trajectories first found by Green in 1839 (Craik, 2004). Inserting the solution so obtained into the particle trajectory ODEs with terms in (29) retained yields another explicit system
@@
Despite the familiarity of the foregoing, a few remarks are in order: the system (32)–(33) which yields the Stokes drift is an approximation of (27)–(28), under the assumption that each particle remains close to its initial position . The result of integrating both systems – (32)–(33) analytically (dashed curves) and (27)–(28) numerically (solid curves) – is shown in Figure 2. Due to the explicit time dependence in the former, the time needed for each particle motion to return to its initial vertical position is exactly equal to the Eulerian period Moreover, after the passage of one Eulerian period each particle has moved an amount in the direction of wave propagation. This drift is indicated by the dashed black curves, which connect the end-points of the dashed trajectories in Figure 2.
By contrast, the time needed for a particle to return to its initial vertical position in the exact, linear theory – which we call the Lagrangian period – is always larger than the Eulerian period. The Lagrangian period is largest at the top of the fluid domain and tends towards the Eulerian period as , as shown also by Constantin & Villari (2008). Moreover, the forward drift of a particle depends on the initial condition, as shown in the black curves connecting the end points of the linear particle trajectories in Figure 2. This drift is sometimes larger and sometimes smaller than the Stokes drift, though its computation using must distinguish between and in the denominator and the corresponding horizontal displacement after a Lagrangian period (for (27)–(28)) or an Eulerian period (for (32)–(33)). Averaging over initial positions covering one wavelength, we find that integration of (27)–(28) yields a larger forward drift than the 2nd order Taylor series approximation (34), as summarised in Figure 3 for several different slopes.
3.2 Nonlinear theory
It is well-known that the deep water potential at third order is formally identical to the first order potential given in (24) (though see Appendix A). Thus, while higher harmonic terms appear in the free surface elevation, the only change to the potential at third order is the nonlinear frequency correction, which means that the characteristic coordinate is now for
In particular, this means that the speed at which each particle moves is faster by an amount of order which impacts the particle trajectories as shown in Figure 4.
For monochromatic waves the procedure used to derive (34) can be applied without alteration to yield the nonlinear corrected Stokes’ drift velocity
Because it is clear that the drift velocity (36) is generally slower than (34) – this despite the fact that nonlinear waves travel faster than their linear counterparts. Moreover, for fixed steepness this correction is more significant for longer waves, i.e. smaller wavenumbers This counterintuitive result – that despite nonlinear waves travelling faster their drift velocity is slower – is due to the fact that the particle trajectories themselves “shrink” with increasing . To lowest order in the Taylor series expansion, where deep-water trajectories are approximately circular, their radius becomes
For the cases considered in Figure 3 the introduction of the frequency correction when leads to a 2 % reduction in the approximate Stokes drift ((36) vs. (34)) and a 5 % reduction in the averaged Lagrangian drift obtained from integrating the particle trajectory mapping with the third-order potential inserted. We will encounter this unexpected consequence of nonlinear theory again in our discussion of more complex wave fields below.

### v2

@@ Abstract
The goal of this work is to investigate particle motions beneath unidirectional, deep-water waves up to the third-order in nonlinearity. A particular focus is on the approximation known as Stokes drift, and how it relates to the particle kinematics as computed directly from the particle trajectory mapping. The reduced Hamiltonian formulation of Zakharov and Krasitskii serves as a convenient tool to separate the effects of weak nonlinearity, in particular the appearance of bound harmonics and the mutual corrections to the wave frequencies. By numerical integration of the particle trajectory mappings we are able to compute motions and resulting drift for sea-states with one, two and several harmonics. We find that the classical formulation provides a slight underestimate of the drift at the surface, and a slight overestimate at depth. Incorporating difference harmonic terms into the formulation yields an improved agreement with the drift obtained from nonlinear wave theories, particularly at greater depth. The consequences of this are explored for regular and irregular waves, as well as parametric wave spectra.
@@ 1 Introduction
Waves on the surface of water are a continuing source of fascination, as well as posing an important scientific and engineering challenge, with impacts on such diverse areas as offshore renewables and marine ecology. While we can observe the myriad forms of waves on the water surface, these shapes are only manifestations of what is going on below the surface: a transport of energy and, to a lesser extent, the movement of water particles themselves. Leonardo da Vinci already observed that the wave outruns the water, and early mathematical analyses of wave kinematics seemed to confirm this idea. Work in the early 19th century by Green showed that particle paths in linear deep water waves are approximately circular – in seeming agreement with the exact, deep-water solution derived by Gerstner – and the corresponding finite depth particle paths were shown to be approximately elliptical by Airy (Craik, 2004).
@@
Nevertheless, much remains unknown from both an experimental and theoretical perspective. Experiments are notoriously difficult, owing to the challenge of comparing a closed flume (with no net mass transport, and with the additional complication of a mechanical wavemaker) with the situation in the (open) ocean. Recent years have seen significant progress, from flume experiments on monochromatic wave trains (Umeyama, 2012; Paprota & Sulisz, 2018) which have been taken to higher and higher steepness (Grue & Kolaas, 2017) to several recent investigations of particle kinematics for wave groups and complex flows (Bjørnestad et al., 2021; van den Bremer et al., 2019).
Meanwhile, theoretical investigations must contend with the fact that the ubiquitous, Eulerian description of fluid motion – dispensing with particle identity in favour of a simpler treatment of velocity fields – necessarily obscures certain aspects of kinematics. While these difficulties can be circumvented for steady flows by the use of alternative formulations, including employing stream function and velocity potential as independent variables as originally suggested by Stokes, there is no easy fix for nonsteady, irregular wave fields. While the pre-eminence of Eulerian theory is being challenged by a resurgence of interest in Lagrangian wave mechanics, including higher-order approximate formulations for steady periodic waves (Clamond, 2007), standing waves (Chen & Hsu, 2009) and, recently, wave packets (Abrashkin & Pelinovsky, 2018; Pizzo et al., 2023), the vast majority of studies continue to employ the Eulerian perspective. The non-trivial task of comparing the two approaches (Fouques & Stansberg, 2008) should not be underestimated.
The aim of this work is to explore the theoretical particle trajectories – and drift – associated with nonlinear potential flow theories in infinite water depth. While there are recent mathematical results for the full, nonlinear problem showing that no closed particle paths exist in steady, periodic wave trains (Constantin, 2006; Henry, 2006; Okamoto & Shoji, 2012), these results are not quantitative and are difficult to generalise to nonsteady problems. We will take a ‘weakly nonlinear’ approach, and our starting point will be the Stokes drift formulation which arises in linear (1st order) water wave theory. While a rigorous proof that linear, steady, periodic water waves exhibit a forward drift in the direction of wave propagation (Constantin & Villari, 2008; Constantin et al., 2008) was recently found, the approximate results due to Stokes are still used in most applications. These were generalised to a spectrum of waves by Kenyon (1969), and the resulting formulation has been extensively explored for different spectral shapes (Webb & Fox-Kemper, 2011), for its sensitivity to the high-frequency spectral tail (Lenain & Pizzo, 2020), and new profiles have been developed for practical implementation (Breivik et al., 2016). Many more references can be found in the recent review by van den Bremer & Breivik (2018).
@@
For a successful comparison we need some way to compute the fluid velocity field , which drives the particle kinematics through the particle trajectory mapping We need a methodology that can deal with several Fourier modes, to simulate irregular seas, and which is accurate to a desired order of nonlinearity. For more than two Fourier modes the Lagrangian theory is currently limited to second order (Pierson, 1961), while important resonant effects in deep water occur only at third order. Consequently, and to increase the accessibility of the results, it seems expedient to work in Eulerian coordinates. Even here several options exist, including the use of traditional perturbation theory to obtain a solution to the water wave problem, which has been extended to third-order by Madsen & Fuhrman (2012). An alternative – which we employ – is to use the equivalent Hamiltonian expansion, which is due to Zakharov (1968) and for which the third-order solution has been explored by Zhang & Chen (1999) and Gao et al. (2021). The Hamiltonian expansion furnishes our velocity field and free surface to third-order, and these are the tools we use to examine
The remainder of this paper is structured as follows: in Section 2 we introduce the water wave problem and corresponding Hamiltonian. We give more details about the expansion that allows us to extract linear, second, and third-order terms in Section 2.1, discuss the neglect of amplitude evolution for our particle kinematics calculations in Section 2.2, and finally show how the free surface and potential are obtained to each order in Section 2.3. In Section 3 we consider monochromatic waves, which is the natural setting to introduce Stokes drift, and look at the consequences of higher-order theories on particle trajectories. Section 4 is devoted to bichromatic waves. These are the simplest type of non-steady wave-field, but because only two harmonics are involved the linear superposition is always periodic. Finally, Section 5 considers waves involving multiple lowest-order harmonics. A discussion of the work and some conclusions are presented in Section 6. Appendix A adds some comments on the non-uniqueness of the monochromatic Stokes wave solution at third-order, while Appendix B gives expressions for some coefficients appearing in Section 2.3. Appendix C explores the effects of slow amplitude evolution on particle paths and Stokes drift.
@@ 2 The water wave problem and Hamiltonian formulation
We begin with the inviscid, irrotational and unidirectional water wave problem in Eulerian variables. This can be written in terms of a fluid velocity potential and free surface as
The advantage of this formulation is that all nonlinearities have been moved to the boundary conditions, which therefore represent the principal challenge. We also note that, provided the water is deeper than about half a typical wavelength, the still-water depth can be taken to be infinite. This procedure, using
@@
in place of (1d) simplifies calculations considerably. Despite the restriction to irrotational flow and the neglect of surface tension, compressibility and viscosity, these equations contain a huge range of interesting physics, including many types of wave phenomena. While the techniques we shall employ can be used without alteration for waves with a transverse () dependence, the additional freedom this entails means that we defer an exploration of directionally spread waves to future work.
During the late 1960s, efforts to find a Hamiltonian formulation of the water wave problem (1a)–(1d) came to fruition in work of Zakharov (1968) who found the Hamiltonian
@@
It is worth noting at the outset that any study of kinematics is necessarily somewhat hampered by the Eulerian variables commonly used in wave research. The Eulerian focus on velocity fields as functions of fixed (laboratory) coordinates obscures the motions of fluid particles, which must be laboriously re-derived. Moreover, as we shall see below, an approximate Eulerian theory divorces the motions of the fluid and the free surface, with consequences for our ability to understand the full kinematics. The alternative choice of Lagrangian variables has many advantages (Pizzo et al., 2023; Blaser et al., 2024), but a general theory of nonlinear, irregular waves has yet to be developed, and the connections to the well-established Eulerian theory remain an topic of active exploration.
@@ 2.2 Constant magnitude approximation
The procedure outlined above would seem to suggest that we first need to solve the complicated integro-differential Zakharov equation (8). In fact, for certain situations of interest we can circumvent this step. The key insight is that the evolution described by the Zakharov equation – being associated with cubic nonlinearities – occurs on the slow timescale , where is a typical wave steepness and is a typical wave period. For processes taking place on faster time-scales, it may be appropriate to consider the complex amplitudes to be constant.
@@
Incorporating these corrections means that the complex amplitudes have constant magnitudes but time-dependent phases
The terms in the argument of the exponential reduce to , the well-known Stokes’ correction, when only a single wave-mode is present, or the mutual frequency correction found by Longuet-Higgins & Phillips (1962) for two modes. The form (10) now contains all the mutual dispersion corrections entailed by the cubically nonlinear water wave theory. It has been demonstrated that taking these corrections into account allows one to predict the short-time propagation of laboratory waves with high accuracy without the need to solve for the evolution of the slow-time variables (Stuhlmeier & Stiassnie, 2021; Galvagno et al., 2021; Meisner et al., 2023), showing that the constant magnitude approximation is well-justified for short times. Of course, under specific circumstances – for example with a sufficiently narrow spectrum, or close to modulation instability – we may have significant energy transfer (Andrade & Stuhlmeier, 2023). However, given our focus on particle paths (with a characteristic time on the order of one period) the amplitude evolution is negligible (see also the discussion in Stuhlmeier & Stiassnie (2019) and Appendix A therein).
If we sum up the consequences of weak-nonlinearity for surface gravity waves in deep water as
1.
dispersion corrections,
2.
changes to wave shape (bound harmonics),
3.
energy exchange between modes,
we will account for 1–2 while neglecting the slow effects 3. Appendix C provides some further discussion of the neglect of amplitude evolution, along with several examples.
@@ 2.3 Recovery of the third-order solution
The corrections contained in the second-order and third-order free surface elevation essentially change the wave shape. The resulting harmonics of the form behave as waves with wavenumber and frequency but because these sums and differences do not satisfy the dispersion relation they are referred to as “bound waves”, in contrast to the free waves of (12). The well-known impact of these bound waves on sharpening the crests and flattening the troughs is shown in Figure 1 for a bichromatic sea, together with the effect of the third-order frequency correction (13) on the wave phases. We also note the important effect of the difference harmonic terms in creating a “set-down” below the crest, see van den Bremer et al. (2019).
@@ 2.3.1 Recovery of the bulk potential
We relate the bulk potential and its surface trace to one another in terms of an expansion. Taking the problem in infinite depth, we need to solve
@@
It should be emphasised that the bulk potential thus obtained is essentially equivalent (except for the non-uniqueness at a given order mentioned in Appendix A) to the potential found through direct perturbation expansion of the governing equations (1a)–(1d’). In particular, as we shall see below, this means that the fluid domain at each order is the half-space owing to the transfer of the boundary conditions.
3 Monochromatic waves
The mathematically simplest type of wave motion, and the one about which we know the most, is that of steady, periodic waves known as Stokes waves. For such waves, symmetric about a crest and propagating without change of form, it can be proven that no closed particle trajectories exist (Constantin, 2006), and explicit calculations of the surface drift have been given by Longuet-Higgins (1987), up to and including the wave of greatest height.
Because the wave-form is steady and the fluid motion is confined to the -plane, a variety of powerful mathematical techniques exist for tackling the problem, including the use of the velocity potential and stream function as independent variables (Stokes, 2009). Moreover, the convergence of perturbative solutions for such waves has been established since the work of Levi-Civita (1925), and in practice such solutions have been computed to extremely high order (Schwartz, 1974). This wealth of prior results means that we do not strictly need the Hamiltonian expansion developed in Section 2; however we shall attempt to situate the 3rd order theory and its predictions among other results for monochromatic waves.
Thus, if we fix the wave length (or wavenumber) and the wave height the successive inclusion of higher order terms brings us ever closer to the exact solution. This appears to be true also when we employ the equivalent Lagrangian formulation of fluid mechanics (Clamond, 2007), albeit without the accompanying analytical theory proving convergence. Some care must be taken in the selection of expansion parameter (in which the “order” is measured), as discussed by Cokelet (1977) and Fenton (1985), about which more will be said below.
Although these steady, periodic waves are theoretically exceptional and experimentally difficult to realise, the wealth of accumulated understanding makes them an ideal starting point, and allows for comparisons that will stand us in good stead for later, unsteady problems. The textbook monochromatic solution (Dean & Dalrymple, 1991) to the linearised version of equations (1a)–(1d’) is
@@ 3 Monochromatic waves
It is worth noting that this system has only a tenuous connection to the linear (or, later, weakly nonlinear) theory, but rather attempts to restore the link between Lagrangian and Eulerian descriptions of the fluid motion. In addition, it is an unwelcome surprise that inserting the linear potential yields a nonlinear system of differential equations
While this system cannot be solved analytically it has been shown rigorously that its particle trajectories are not closed (Constantin & Villari, 2008). The result is qualitative, and an analytical treatment that yields the drift quantitatively even in this simple case appears out of reach.
In lieu of this, two alternatives remain: approximation or numerical solution. The former approach is found in most textbooks on water waves, and begins with a Taylor expansion of the fluid velocity field about an initial position . Keeping only the lowest order terms in the Taylor expansion yields, upon integration, the circular particle trajectories first found by Green in 1839 (Craik, 2004). Inserting that solution into the second-order Taylor expansion yields another explicit system of ODEs
@@
Figure 2 compares the solution of the linear particle trajectory mapping (28)–(29) with that of its approximation (30)–(31). Following Longuet-Higgins (1987), we define the Lagrangian period as
where is the Eulerian period, is the phase speed of the wave, and is the Stokes drift at a given depth. In the approximate model, shown with dashed curves in Figure 2, after one Eulerian period each particle has moved a uniform amount to the right (diamond markers). By contrast, after one Lagrangian period, calculated using (32), the linear theory yields a forward drift that depends on the initial position (square markers), and which is sometimes smaller and sometimes larger than the Stokes drift. This dependence of the drift on the phase was also observed experimentally by Grue & Kolaas (2020).
More importantly, Figure 2 emphasises a crucial blind-spot of the linear theory: the fluid velocity field is defined only below Thus even a particle starting initially at trough level – such as those on the right side of Figure 2 – will enter a region of the -plane where the velocities are undefined (indeed, they may also exceed the crest height, as can be seen in the red trajectory on the far right; we should not make too much of this fact, since the free surface in linear (Eulerian) theory is not a streamline nor is it composed of fluid particles).
An evaluation of the velocity field above amounts to an extrapolation of the linear theory, a procedure which is known to overestimate velocities near the surface (Johannessen, 2010). One possibility to obtain surface drift values is integration of the horizontal surface velocity (see Grue & Kolaas (2020), who use this to 2nd order to compute the Lagrangian period) where the right-hand side can be approximated as
As mentioned, we do not strictly require the Hamiltonian expansion for monochromatic waves, and in practice we make use of asymptotic expansions up to fifth order in which were derived by Stokes for deep water and are reviewed in the comprehensive recent paper by Zhao & Liu (2022).
The results are shown in Figure 3, which compares the formula (32) (label SD) with weakly nonlinear approximations up to 5th order (omitting the 1st order, which is effectively unusable), as well as the exact surface drift obtained by Longuet-Higgins (1987) (label LH). Taking 1/m, the horizontal axis denotes half the actual crest-to-trough height, and not the leading-order amplitude which is used in the perturbation expansion, and which must be adjusted every odd order to obtain a desired value of (see discussion in Appendix A). Without such an adjustment the wave height grows with the order of the expansion, as does the surface drift.
Figure 3 shows that for waves of low steepness (below ) the differences in the formulations are essentially negligible. The Stokes drift formula (32) gives a drift that is slightly too small (as shown by Longuet-Higgins (1987) and complementary work using the Lagrangian formulation, such as Clamond (2007)), but higher-order theories provide good agreement up to very high steepness (though the trend clearly does not continue to the steepest wave with ). We note that dispersion corrections appearing at the 3rd and 5th order have the effect of decreasing surface drift slightly, while the seemingly excellent agreement between the 2nd order approximation and the exact solution at slopes above should be considered essentially accidental.
The depth-dependent Stokes drift can be obtained from integrating the particle trajectory mapping, provided particles do not cross the still-water level – otherwise, as mentioned above, extrapolating the terms can lead to spuriously high velocities or even divergence of the trajectories. Because the forward drift of a particle is phase-dependent, this means averaging over the phases of particles lying at an initial level In Figure 4, 50 points covering the wave phase are used at each depth to compute the effective Stokes drift. We compare results obtained from 1st, 3rd and 5th order velocity fields, as well as a 4th order Stokes drift from the Lagrangian obtained by Blaser et al. (2025) (see also Clamond (2007)), and which we shall return to for unsteady cases below.
It is noteworthy that the integration of the 1st order velocity field (red curves, label 1) gives a significant overestimate of the Lagrangian drift with depth, while the approximation (32) based on that same velocity field (blue curves, label SD) gives much better agreement with both the 4th order Lagrangian result and higher-order Eulerian velocity fields. Only for high steepness and shown in panels (b)–(c) is the approximate Stokes drift noticeably different from the higher-order solutions, although its asymptotics at large depth hew close to the 1st order theory, while the higher-order solutions have a somewhat different behaviour at large depth (see inset in panel (c)).
In the formulation of the velocity potential we are using here (corresponding to (44a)) the only difference in velocity potential between the 1st and 3rd order is only the inclusion of nonlinear dispersion, so that waves travel with a characteristic coordinate for
No additional harmonics appear until 4th order, so the difference between curves 1 and 3 in Figure 4 is entirely a consequence of this dispersion correction. Indeed, the procedure used to derive (32) can be applied without alteration to the 3rd order potential, yielding
Because it is clear that the drift velocity (35) is generally smaller than (32), as observed also from the particle trajectories.
We can summarise the results for monochromatic waves as follows: the approximate Stokes drift formula (32) gives results that are somewhat too small at the surface, and somewhat too large at depth. The differences for waves of moderate steepness, however, are not sizeable, as borne out also by experimental work (Paprota & Sulisz, 2018). In our description, the only effects of nonlinearity are in adjusting the frequency (which has consequences for the entire flow) and in the appearance of higher-harmonics (whose effect is confined close to the surface due to the -terms). We shall now see how these effects appear in unsteady (bichromatic) wave trains.